OpenStudy (anonymous):

Reina and Sam are multiplying (84)5 (73)9. @zaibali.qasmi

OpenStudy (anonymous):

Reina's Work Sam's Work (84)5 (73)9 = 84 + 573 + 9 = 89712 (84)5 (73)9 = 84⋅573⋅9 = 820727' Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

all are multiplying..?

OpenStudy (anonymous):

i dont get your question..

OpenStudy (anonymous):

outside the parenthesis your multiplying right?

OpenStudy (anonymous):